OpenStudy (anonymous):

A $6000 investment earns $690 simple interest in two years. What is the annual interest rate? A. 1.15% B. 5.75% C. 11.5% D. 57.5% I NEED HELP!!! DESPERATLY!!! OH PLEASE PLEASE PLEASE PLEASE PLEASE!!!! I WILL GIVE MEDAL IF SOMEONE HELP ME!!

OpenStudy (anonymous):

@ribhu @ogjb_ @e.mccormick @TwiztTiez @chainedecho @TheEdwardsFamily @flasht @kathlyn98 @LeeEtchison @kathlyn98 @Jamielynne @ubiquitous @fallenangelorchid @chainedecho

OpenStudy (anonymous):

@AdamK plz help

OpenStudy (adamk):

11.5%

OpenStudy (anonymous):

may u plz explain how?

OpenStudy (anonymous):

Please stop tagging me, I thought I said that I'm not very good at this type of problem.

OpenStudy (anonymous):

wow, your good, almost too good

OpenStudy (anonymous):

u did not

OpenStudy (anonymous):

at least i did not see it

OpenStudy (anonymous):

orry :(

OpenStudy (anonymous):

sorry

OpenStudy (adamk):

Here, it's the rate that is unknown, so you solve for r: I = Prt r = I/Pt = 690/(6000*1) = 0.115 = 11.5%

OpenStudy (anonymous):

Oh sorry I was thinking of another person--my bad XD

OpenStudy (pinklion23):

Hey it is B i have the same question on my test and it was B.